Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Siblings

Expert replies
by Jinglander » Tue Aug 10, 2010 7:50 pm
In a room filled with 7 people, 4 have 1 sibling and 3 have two siblings. If two people at selected at random what is the prob that they are not siblings.

Answer is 6/21

Can someone explain.
Join the discussion
Source: — Problem Solving |

by Rahul@gurome » Tue Aug 10, 2010 8:49 pm
Jinglander wrote:In a room filled with 7 people, 4 have 1 sibling and 3 have two siblings. If two people at selected at random what is the prob that they are not siblings.

Answer is 6/21

Can someone explain.
4 people have 1 sibling means there are 2 pairs of siblings.
Let A and B be the siblings, and C and D be the other pair of siblings.

Since 3 people have 2 siblings, so let E, F and G be the pair of 3 siblings.

Total no. of favorable ways of selecting 2 people from 7 people = 7C2 = 7!/(2!)(5!) = 21 ways
No. of favorable ways of selecting a sibling pair = 1 + 1 + 3 = 5 ways (We can select pairs from A-B, C-D, E-F, F-G, E-G)

Probability that they are siblings = 5/21
Probability that they not siblings = 1 - 5/21 = 16/21
Rahul Lakhani
Quant Expert
Gurome, Inc.
https://www.GuroMe.com
On MBA sabbatical (at ISB) for 2011-12 - will stay active as time permits
1-800-566-4043 (USA)
+91-99201 32411 (India)
Join the discussion

by sheelanadh » Wed Aug 11, 2010 4:24 am
Rahul@gurome wrote:
Jinglander wrote:In a room filled with 7 people, 4 have 1 sibling and 3 have two siblings. If two people at selected at random what is the prob that they are not siblings.

Answer is 6/21

Can someone explain.
4 people have 1 sibling means there are 2 pairs of siblings.
Let A and B be the siblings, and C and D be the other pair of siblings.

Since 3 people have 2 siblings, so let E, F and G be the pair of 3 siblings.

Total no. of favorable ways of selecting 2 people from 7 people = 7C2 = 7!/(2!)(5!) = 21 ways
No. of favorable ways of selecting a sibling pair = 1 + 1 + 3 = 5 ways (We can select pairs from A-B, C-D, E-F, F-G, E-G)

Probability that they are siblings = 5/21
Probability that they not siblings = 1 - 5/21 = 16/21
thank you....is the way of approach is correct..what is correct OA?
Join the discussion